# Superiorization and Perturbation Resilience of Algorithms: A Continuously Updated Bibliography

### v9

@@ Abstract
Since the beginnings of this topic we try to trace the work that has been published about it since its inception. To the best of our knowledge this bibliography represents all available publications on this topic to date, and while the URL is continuously updated we will revise this document and bring it up to date on arXiv approximately once a year. Abstracts of the cited works, and some links and downloadable files of preprints or reprints are available on the above mentioned Internet page. If you know of a related scientific work in any form that should be included here kindly write to me on: yair@math.haifa.ac.il with full bibliographic details, a DOI if available, and a PDF copy of the work if possible. The Internet page was initiated on March 7, 2015, and has been last updated on March 9, 2024.

### v10

@@ Abstract
Since the beginnings of this topic we try to trace the work that has been published about it since its inception. To the best of our knowledge this bibliography represents all available publications on this topic to date, and while the URL is continuously updated we will revise this document and bring it up to date on arXiv approximately once a year. Abstracts of the cited works, and some links and downloadable files of preprints or reprints are available on the above mentioned Internet page. If you know of a related scientific work in any form that should be included here kindly write to me on: yair@math.haifa.ac.il with full bibliographic details, a DOI if available, and a PDF copy of the work if possible. The Internet page was initiated on March 7, 2015, and has been last updated on March 19, 2025.
@@ 2 The Bibliography
[179] Alexander J. Zaslavski, Solutions of Fixed Point Problems with Computational Errors, Springer Nature Switzerland AG, (2024).
[180] Alexander J. Zaslavski, Convergence of inexact orbits of nonexpansive mappings in complete metric spaces, Communications in Optimization Theory, Vol. 2024 (2024), Article ID 4, pp. 1–10.
https://cot.mathres.org/issues/COT20244.pdf.
[181] Steve Collins, Andy Ogilvy, Warren Hare, Michelle Hilts and Andrew Jirasek, Iterative Image Reconstruction Algorithm Analysis for Optical CT Radiochromic Gel Dosimetry, analysis, and performance evaluation, Biomedical Physics & Engineering Express, accepted manuscript, open access (2024). https://iopscience.iop.org/article/10.1088/2057-1976/ad3afe.
[182] Rodrigo Hernangómez, Jochen Fink, Renato L.G. Cavalcante, Zoran Utkovski and Slawomir Stańczak, Optimized Detection with Analog Beamforming for Monostatic Integrated Sensing and Communication, 2024. Available on arXiv at: https://arxiv.org/abs/2404.08455.
[***] F.J. Aragón-Artacho, Y. Censor, A. Gibali and D. Torregrosa-Belén, Corrigendum to: The superiorization method with restarted perturbations for split minimization problems with an application to radiotherapy treatment planning, Applied Mathematics and Computation, Vol. 476, Article 128788 (2024). Available at:
https://www.sciencedirect.com/science/article/pii/S0096300324002522.
[183] F.J. Aragón-Artacho, W. Cai, Y. Censor, A. Gibali, C. Shui and D. Torregrosa-Belén, Approaches to iterative algorithms for solving nonlinear equations with an application in tomographic absorption spectroscopy, Communications in Optimization Theory, accepted for publication, 2024. Available on arXiv at: https://arxiv.org/abs/2405.08635.
[184] Shousheng Luo, Zhiting Liu, Yaofei Lu and Xue-Cheng Tai, Superiorized iteration algorithm for CT image simultaneous reconstruction and segmentation, Inverse Problems and Imaging, DOI:10.3934/ipi.2024024, (2024). Available at American Institute of Mathematical Sciences.
https://www.aimsciences.org//article/doi/10.3934/ipi.2024024.
[185] S. Yu, Y. Censor and G. Luo, Floorplanning With I/O Assignment via Feasibility-Seeking and Superiorization Methods, IEEE Transactions on Computer-Aided Design of Integrated Circuits and Systems, Vol. 44, pp. 317–330, (2024). Online at:
IEEEXplore. https://ieeexplore.ieee.org/abstract/document/10543062.
[186] David Torregrosa Belen, Splitting Algorithms for Structured Optimization: Theory and Applications, Ph.D. Dissertation, Department of Mathematics, University of Murcia, Spain, 2024. Dissertation available at: Digitum: Repositorio Institucional de la Universidad de Murcia.
https://digitum.um.es/digitum/handle/10201/142204.
[187] Touraj Nikazad, Perturbed conjugate gradient method: controlling semi-convergence, enforcing non-negativity, and accelerating iterations in tomographic imaging, Physica Scripta, Vol. 99, 075034, (2024). DOI:10.1088/1402-4896/ad5651. Available at: IOPscience.
https://iopscience.iop.org/article/10.1088/1402-4896/ad5651/meta.
[188] Jan Schröder, Yair Censor, Philipp Süss and Karl-Heinz Küfer, Immunity to increasing condition numbers of linear superiorization versus linear programming, Preprint. Available on arXiv at: https://arxiv.org/abs/2407.18709.
[189] Chen JingRuo, Xu ShiJie, Liu HeCong, Huang JianQing, Liu YingZheng and Cai WeiWei, Untrained neural network for linear tomographic absorption spectroscopy, Science China Technological Sciences, Vol. 67, (2024). https://doi.org/10.1007/s11431-023-2629-2. Available on SpringerLink.
[190] Jon Henshaw, Aviv Gibali and Thomas Humphries, Plug-and-play superiorization. Available on arXiv at: https://arxiv.org/abs/2410.23401.
[191] Satoshi Shoji, Wataru Yata, Keita Kume and Isao Yamada, A Discrete-Valued Signal Estimation by Nonconvex Enhancement of SOAV with cLiGME Model, 2024 Asia Pacific Signal and Information Processing Association Annual Summit and Conference (APSIPA ASC), GALAXY INTERNATIONAL CONVENTION CENTER, Macau Cotai Macau, Macau (2024). http://www.apsipa2024.org/welcome.html.
[192] K. Barshad, Y. Censor, W.M. Moursi, T. Weames and H. Wolkowicz, A necessary condition for the guarantee of the superiorization method, Optimization Letters, https://doi.org/10.1007/s11590-025-02192-7 (2025). Open Access at Springer Nature Link.
[193] Satoshi Shoji, Wataru Yata, Keita Kume and Isao Yamada, An LiGME Regularizer of Designated Isolated Minimizers – An Application to Discrete-Valued Signal Estimation, (2025). Available on arXiv at:
https://arxiv.org/abs/2503.10126.